Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 186     Entry time: 06 May 2023, 06:53
Author: Giuseppe Mercurio 
Group: SCS 
Subject: night shift summary 

achievements:

- measured delay scans at sample tilted by 20 degree

problems:

- we saw a drop in pulse energy around 00:30 in XGM at SCS, we verified that it was not because of the FEL not going through the HSLITS, in the end after calling BKR it was realized that there was indeed a problem with BC2, after fixing this, beam pointing was recovered and pulse energy as well

- beam pointing instabilities during tuning for SA1 around 6:30

ELOG V3.1.4-7c3fd00